Grade 12Physical Chemistry

At 25degre Celsius the vapour pressure of pure water is 23.76 mmHg and tgat of aquesous dilute solution of urea is 22.98mm of hg what is the molarity of the solution

Chhaya
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
 
Dear Chhaya
 
we know that,
total vapour pressure = pure vapour pressure X mole fraction
22.98  =  23.76 X Xw
therefore,
Xw = 22.98/23.76
       =Nw/Nw+ Na
so,
22.98/23.76--22.98
====>Nw/Na
therefore,
Na/Nw = 23.76--22.98/22.98
 
NOW,
molality = Na/mass of water in kg
             =Na/Nw X 18X10^-3
 so molality = 23.76--22.98/22.98 X 18X10^-3
so the final answer = 0.001X 10^-3
 
 
Regards
Arun (askIITians forum expert)